OpenStudy (anonymous):

Solve log5 ( x + 2) + log5 ( 2x - 1) = 2

jimthompson5910 (jim_thompson5910):

log5 ( x + 2) + log5 ( 2x - 1) = 2 log5[ ( x + 2)*( 2x - 1) ] = 2 ( x + 2)*( 2x - 1) = 5^2 2x^2 - x + 4x - 2 = 25 2x^2 + 3x - 2 = 25 I'll let you finish. Don't forget to check any possible answers you get.